OpenStudy (anonymous):

find the invers of the one-to-one function f(x)=4/5x-1

OpenStudy (anonymous):

answer key says it is f^-1(x)=4/5x+1/5

OpenStudy (idku):

\[y=\frac{4}{5x-1}\]\[x=\frac{4}{5y-1}\]\[x(5y-1)=4\]\[5y-1=\frac{4}{x}\]\[5y=\frac{4}{x}+1\]\[5y=\frac{4}{x}+\frac{x}{x}\]\[5y=\frac{4+x}{x}\]\[y=\frac{4+x}{5x}\]

OpenStudy (idku):

\[f^{-1}(x)=\frac{x+4}{5x}\]

OpenStudy (anonymous):

answer key says4/5x+1/5 ????

OpenStudy (anonymous):

is the answer key wrong ?????????????????????????????????????????????????

OpenStudy (anonymous):

I also got 4+x/5x

OpenStudy (anonymous):

this is my test review im calm im just thinking maybe the teacher did a typo my answer key to this problem is 4/5x +1/5

OpenStudy (kmeis002):

Note that: \[ \frac{4}{5x} + |\frac{1}{5} = \frac{4}{5x} + \frac{x}{5x} = \frac{4+x}{5x} = \frac{x+4}{5x}\] Note that these answers are the same, just in different form

OpenStudy (anonymous):

I see that thanks